# How Far are VLMs from Visual Spatial Intelligence? A Benchmark-Driven Perspective

### v1

@@ 6.1 Challenges
Lack of Precise and Quantitative Capabilities. VLMs exhibit inherent shortcomings in precision and quantification, a limitation that may stem from their reliance on learning from human language, which is often ambiguous when describing the physical world. Consequently, while these models excel at processing qualitative spatial relationships, such as ”A is to the left of B,” they demonstrate significant deficiencies in tasks requiring quantitative geometric computation. This deficiency is empirically validated by benchmark results. As illustrated in Fig. 1, existing models consistently underperform on tasks that demand precise numerical regression, including Coordinate Conversion, Distance Prediction, and Object Size Estimation. This difficulty in mapping visual features to precise physical coordinates or metric units severely restricts the application potential of VLMs in high-precision scenarios, such as robotic grasping and the accurate placement of content in Augmented Reality (AR).
@@
Insufficiency in Dynamic-Temporal and Cross-View Reasoning. The real world is inherently dynamic and variable, necessitating that an intelligent agent integrate information across time and viewpoints to form a coherent spatial cognition. However, current VLMs demonstrate notable insufficiencies in this area. The Fig. 1 highlights the models’ weaknesses in tasks such as Multi-View Reasoning, Velocity & Acceleration, and Trajectory Description. This suggests that VLMs struggle both to track the spatio-temporal state of objects effectively and to align and fuse information from disparate camera viewpoints. This deficiency in processing dynamic and multi-view information poses a critical barrier to the advancement of VLMs into high-level applications requiring continuous interaction with dynamic environments, such as robotics and autonomous driving.
@@ 6.3 Applications
As this review has elaborated, VSR represents not only a theoretical challenge but also the critical capability for translating digital intelligence into real-world action. The profound impact of VSR is most vividly demonstrated across different frontier domains, such as autonomous driving, embodied robotics, and augmented reality.
Embodied Intelligence. For embodied AI and robotics, VSR acts as the crucial bridge connecting abstract instructions to concrete physical actions [31, 227]. It enables agents to ground natural language commands (e.g.“place the cup in the center of the table”) into precise 3D spatial coordinates and action sequences [214], thereby facilitating complex navigation [31, 33] and manipulation tasks [215]. Although current models still exhibit deficiencies in geometric and physical commonsense reasoning, the rise of Vision-Language-Action (VLA) models [228, 229], particularly those incorporating internal reasoning mechanisms like the “visual chain-of-thought” [230], is advancing robotic intelligence toward a ”think before you act” paradigm. This approach effectively decomposes complex tasks and improves the success rate of planning and execution.
Autonomous Driving. In the domain of autonomous driving, VSR serves as the cognitive cornerstone for safe navigation. It elevates a vehicle’s perceptual capabilities from simple object recognition to deep understanding of the 3D structure of dynamic traffic scenes [231, 232, 233], the relative relationships between objects [234], and their motion trajectories [235]. This capability is a prerequisite for all advanced planning and decision-making. A central challenge in this field is bridging the gap between the 2D image-based reasoning abilities developed during model training and the need for precise, quantitative inference in the 3D physical world [236]. To this end, the field is shifting from imitation learning towards explicit reasoning, developing end-to-end models like EMMA [117] and neuro-symbolic hybrid frameworks such as Logic-RAG [237], to build more robust and interpretable driving intelligence.
@@ 7 Conclusion
In this paper, we presents a systematic and comprehensive investigation into the domain of Visual Spatial Reasoning (VSR) for Vision-Language Models (VLMs). We begin by conducting a review of existing methodologies, analyzing them through input modalities, model architectures, training strategies, and inference mechanisms. Concurrently, we introduce a cognition-based hierarchical taxonomy that categorizes VSR tasks into three distinct levels: Basic Perception, Spatial Understanding, and Spatial Planning. Building upon this framework, we curate SIBench, a comprehensive benchmark that integrates nearly 20 open-source datasets across 23 task settings to facilitate a holistic evaluation of VLM capabilities. Our experiments on state-of-the-art models using SIBench reveal a critical gap between perceptual and cognitive abilities. While models demonstrate proficiency in basic perception, they exhibit significant deficiencies in higher-order reasoning, particularly in tasks requiring precise numerical estimation, multi-view reasoning, dynamic temporal understanding, and spatial imagination. These findings underscore the substantial challenges that remain on the path to achieving spatial intelligence in artificial agents.
@@ Appendix A Appendix Outline
This supplementary material includes 5 aspects:
@@ A.3 Detailed Comparison of Various Settings
SIBench is a comprehensive benchmark consisting of 23 distinct task settings. The detailed results of various models on SIBench and SIBench-mini are shown in Table V and Table VI, respectively. Models generally excel at basic perception tasks like identifying the existence of objects but struggle significantly with tasks requiring complex, abstract reasoning, such as spatial imagination, temporal ordering, and precise quantitative estimation. Overall, Gemini-2.5-Pro [21] is the top one in the SIBench, while the GPT-5 [20] leads in the SIBench-mini.
@@ A.3.1 Areas of Model Strength
In Table V and Table VI, models consistently perform well on tasks that rely on recognizing objects and simple relationships of objects. On the existence task, Qwen2.5-VL-72B [41] achieves a perfect score of 1.0, and many others are above 0.9. This indicates that identifying whether an object is present in a scene is an easy task for current models. On the Spatial Relation task, InternVL-3-78B [42] leads at 0.8402 and Qwen2.5-VL-72B [41] also receives 0.8216. This strength suggests that models have a solid grasp of static relational ability.
@@ A.3.2 Areas of Model Weakness
The models exhibit several key weaknesses, consistent with the analysis in Section 5.2.2. First, models are weak at some tasks of Foundational Perception, such as Object Size Estimation shown in Table V and Table VI. Second, models struggle with quantitative reasoning. This is evident in tasks Relative Distance in Table V and Table VI. The scores indicate that models like Gemini-2.5-Flash [21] and InternVL-3-78B [42] rely on coarse visual clues rather than precise metric representations. Third, models show a near-total absence of spatial imagination. On the Spatial Imagination task, models like GPT-4o-mini [224] and InternVL-2.5-78B-MPO [42] receive almost the lowest scores. This indicates that models fail on tasks that require the creation and manipulation of a mental imagination. Models cannot visualize and reason about objects or scenarios that are not explicitly depicted, which is a key component of human-like spatial intelligence. Second, models are weak at inferring spatial properties from different viewpoints. This is shown by the low scores in Multi-View Reasoning, Camera Pose, and Geometric Reasoning. Models like Qwen2.5-VL-7B [41], LLaVA-OneVision-72B [18], Gemini-2.5-Flash [21] cannot intuitively understand how a camera’s movement changes a scene’s perspective, leading to poor performance. Similarly, models struggle with temporal appearance order, with LLaVA-OneVision-7B [18] scoring 0.0000. This confirms that models’ unreliable ability for non-static scenes.
TABLE V: Performance evaluation of different models on visual spatial reasoning tasks. Bold is the best, and underlining is the second best.
@@
Basic Perception Reach Prediction $0.6000$ 0.6750 0.6750 0.6000 0.6500 0.4750 0.5750 0.6250 0.5750 0.5500
Height $0.5117$ 0.6033 0.6117 0.6400 0.5883 0.6167 0.6050 0.6233 0.5733 0.7133
Existence 0.9750 1.0000 0.9000 0.9500 0.9250 0.9750 0.9250 0.9500 0.9000 0.9000
Occlusion $0.5660$ 0.5600 0.6620 0.6640 0.5880 0.6160 0.7160 0.7980 0.6160 0.7300
Object Shape $0.2231$ 0.3462 0.2846 0.3077 0.2692 0.3692 0.3308 0.3231 0.3000 0.4154
Counting 0.5588 0.5504 0.5042 0.5084 0.4454 0.3991 0.5126 0.4495 0.4916 0.4664
Object Size Estimation $0.4966$ 0.6085 0.5459 0.6045 0.5530 0.4575 0.5590 0.4573 0.6139 0.5968
Spatial Understanding Spatial Compatibility $0.5280$ 0.3925 0.4673 0.4019 0.4813 0.4673 0.5701 0.5841 0.5748 0.7103
Coordinate Conversion $0.5100$ 0.6525 0.5413 0.5950 0.5113 0.3563 0.5950 0.6450 0.5900 0.6900
Trajectory Description $0.3205$ 0.6026 0.3372 0.5513 0.2179 0.4615 0.3718 0.2692 0.5513 0.3333
Geometric Reasoning $0.2103$ 0.2421 0.2738 0.2579 0.2897 0.1786 0.2778 0.2971 0.2659 0.3730
Spatial Imagination 0.3100 0.2950 0.3075 0.2650 0.2325 0.2575 0.2425 0.2700 0.3100 0.2700
Spatial Grid $0.4675$ 0.8050 0.4025 0.5200 0.4375 0.5650 0.6375 0.6475 0.7525 0.9975
Temporal Appearance Order $0.1818$ 0.4545 0.0000 0.3636 0.2727 0.3636 0.1818 0.0909 0.2727 0.5455
Multi-View Reasoning $0.1718$ 0.4125 0.3892 0.4295 0.3287 0.4358 0.3775 0.3637 0.3998 0.4687
Situational QA $0.4961$ 0.5295 0.4972 0.5095 0.4729 0.4036 0.4838 0.5217 0.5173 0.5474
Velocity Acceleration $0.3613$ 0.4622 0.3882 0.4403 0.2605 0.3344 0.4202 0.4840 0.5092 0.4034
Relative Distance $0.2920$ 0.5438 0.4629 0.4464 0.3130 0.3608 0.3645 0.4189 0.5462 0.5459
Camera Pose $0.4041$ 0.4306 0.3372 0.2871 0.3198 0.3459 0.5029 0.5581 0.5087 0.4041
Spatial Relation 0.8216 0.8029 0.7448 0.7884 0.5892 0.4395 0.7905 0.8402 0.8133 0.6660
Object Localization $0.1994$ 0.3769 0.7290 0.7570 0.5763 0.5670 0.2523 0.3832 0.2461 0.7041
Planning Maze Navigation $0.2775$ 0.4350 0.3375 0.3975 0.2950 0.6400 0.5725 0.4950 0.4850 0.8625
Route Planning $0.3220$ 0.2881 0.3220 0.3220 0.3729 0.4068 0.3390 0.2542 0.4576 0.3898
Overall $0.4172$ 0.5114 0.4844 0.5103 0.4278 0.4389 0.4983 0.5197 0.5252 0.5883
TABLE VI: Performance evaluation of different models on the mini dataset. Bold is the best, and underlining is the second best.
@@
Basic Perception Reach Prediction $0.5750$ 0.6000 0.5750 0.6500 0.5000 0.5750 0.6000
Height $0.5362$ 0.6522 0.5633 0.7750 0.7500 0.7500 0.8000
Existence $0.8750$ 0.9500 0.8750 0.9750 0.9500 0.9750 0.9250
Occlusion $0.6250$ 0.6000 0.6160 0.7250 0.8000 0.7250 0.6250
Object Shape $0.3000$ 0.2500 0.3000 0.4000 0.3250 0.2000 0.4750
Counting $0.8333$ 0.8889 0.4916 0.7222 0.9444 0.8333 0.7222
Object Size Estimation $0.5150$ 0.5075 0.6139 0.5800 0.5450 0.6175 0.6425
Spatial Understanding Spatial Compatibility $0.6250$ 0.7143 0.5748 0.7000 0.6500 0.7250 0.6250
Coordinate Conversion $0.6250$ 0.5500 0.5900 0.8000 0.7250 0.7000 0.6750
Trajectory Description 0.5500 0.5500 0.5513 0.3750 0.4500 0.4500 0.4750
Geometric Reasoning $0.3000$ 0.2750 0.2659 0.3000 0.3000 0.4000 0.3500
Spatial Imagination $0.2500$ 0.3750 0.3100 0.3250 0.3250 0.3250 0.3250
Spatial Grid $0.7250$ 0.6250 0.7265 0.9250 0.9750 1.0000 0.7750
Temporal Appearance Order $0.1818$ 0.3636 0.2727 0.1818 0.9091 0.1818 1.0000
Multi-View Reasoning $0.2250$ 0.3750 0.2269 0.3750 0.5250 0.4500 0.5750
Situational QA $0.3438$ 0.3750 0.5173 0.3438 0.5000 0.4375 0.6250
Velocity Acceleration $0.4500$ 0.4000 0.5092 0.4750 0.5750 0.5000 0.5250
Relative Distance $0.4675$ 0.4393 0.4811 0.6775 0.7350 0.6675 0.7075
Camera Pose 0.5500 0.2292 0.4884 0.3000 0.4000 0.5000 0.4000
Spatial Relation $0.7500$ 0.7750 0.8133 0.7500 0.6500 0.6750 0.6000
Object Localization $0.2326$ 0.4651 0.2430 0.7500 0.8250 0.7750 0.8000
Planning Maze Navigation $0.3571$ 0.3810 0.3150 0.7750 0.8500 0.7750 0.7500
Route Planning 0.5000 0.3750 0.4576 0.2500 0.4500 0.3750 0.4500
Overall $0.4987$ 0.5006 0.5355 0.5822 0.6216 0.6295 0.6906

### v2

@@ 6.1 Challenges
Lack of Precise and Quantitative Capabilities. VLMs exhibit inherent shortcomings in precision and quantification, a limitation that may stem from their reliance on learning from human language, which is often ambiguous when describing the physical world. Consequently, while these models excel at processing qualitative spatial relationships, such as ”A is to the left of B,” they exhibit significant deficiencies in tasks requiring quantitative geometric computation. This deficiency is empirically validated by benchmark results. As illustrated in Fig. 1, existing models consistently underperform on tasks that demand precise numerical regression, including Coordinate Conversion, Distance Prediction, and Object Size Estimation. This difficulty in mapping visual features to precise physical coordinates or metric units severely restricts the application potential of VLMs in high-precision scenarios, such as robotic grasping and the accurate placement of content in Augmented Reality (AR).
@@
Insufficiency in Dynamic-Temporal and Cross-View Reasoning. The real world is inherently dynamic and variable, necessitating that an intelligent agent integrate information across time and viewpoints to form a coherent spatial cognition. However, current VLMs exhibit notable insufficiencies in this area. The Fig. 1 highlights the models’ weaknesses in tasks such as Multi-View Reasoning, Velocity & Acceleration, and Trajectory Description. This suggests that VLMs struggle both to track the spatio-temporal state of objects effectively and to align and fuse information from disparate camera viewpoints. This deficiency in processing dynamic and multi-view information poses a critical barrier to the advancement of VLMs into high-level applications requiring continuous interaction with dynamic environments, such as robotics and autonomous driving.
@@ 6.3 Applications
As this review has elaborated, VSR represents not only a theoretical challenge but also the critical capability for translating digital intelligence into real-world action. The profound impact of VSR is most vividly exhibit across different frontier domains, such as autonomous driving, embodied robotics, and augmented reality.
Embodied Intelligence. For embodied AI and robotics, VSR acts as the crucial bridge connecting abstract instructions to concrete physical actions [VSI-Bench, krantz_vlnce_2020]. It enables agents to ground natural language commands (e.g.“place the cup in the center of the table”) into precise 3D spatial coordinates and action sequences [wang2024vision], thereby facilitating complex navigation [VSI-Bench, OmniSpatial] and manipulation tasks [gong2025space]. Although current models still exhibit deficiencies in geometric and physical commonsense reasoning, the rise of Vision-Language-Action (VLA) models [qu2025spatialvla, zheng2024tracevla], particularly those incorporating internal reasoning mechanisms like the “visual chain-of-thought” [Cot-VLA], is advancing robotic intelligence toward a ”think before you act” paradigm. This approach effectively decomposes complex tasks and improves the success rate of planning and execution.
Autonomous Driving. In the domain of autonomous driving, VSR serves as the cognitive cornerstone for safe navigation. It elevates a vehicle’s perceptual capabilities from simple object recognition to deep understanding of the 3D structure of dynamic traffic scenes [zhou2024embodied, cao2024maplm, tian2025nuscenes], the relative relationships between objects [choudhary2024talk2bev], and their motion trajectories [sachdeva2024rank2tell]. This capability is a prerequisite for all advanced planning and decision-making. A central challenge in this field is bridging the gap between the 2D image-based reasoning abilities developed during model training and the need for precise, quantitative inference in the 3D physical world [zhou_vlmad_survey]. To this end, the field is shifting from imitation learning towards explicit reasoning, developing end-to-end models like EMMA [hwang2024emma] and neuro-symbolic hybrid frameworks such as Logic-RAG [kabir2025logic], to build more robust and interpretable driving intelligence.
@@ 7 Conclusion
In this paper, we presents a systematic and comprehensive investigation into the domain of Visual Spatial Reasoning (VSR) for Vision-Language Models (VLMs). We begin by conducting a review of existing methodologies, analyzing them through input modalities, model architectures, training strategies, and inference mechanisms. Concurrently, we introduce a cognition-based hierarchical taxonomy that categorizes VSR tasks into three distinct levels: Basic Perception, Spatial Understanding, and Spatial Planning. Building upon this framework, we curate SIBench, a comprehensive benchmark that integrates nearly 20 open-source datasets across 23 task settings to facilitate a holistic evaluation of VLM capabilities. Our experiments on state-of-the-art models using SIBench reveal a critical gap between perceptual and cognitive abilities. While models exhibit proficiency in basic perception, they exhibit significant deficiencies in higher-order reasoning, particularly in tasks requiring precise numerical estimation, multi-view reasoning, dynamic temporal understanding, and spatial imagination. These findings underscore the substantial challenges that remain on the path to achieving spatial intelligence in artificial agents.
@@ Appendix A Appendix Outline
This supplementary material includes 6 aspects:
@@ A.3 Detailed Comparison of Various Settings
SIBench is a comprehensive benchmark consisting of 23 distinct task settings. The detailed results of various models on SIBench and SIBench-mini are shown in Table IV and Table V, respectively. Models generally excel at basic perception tasks like identifying the existence of objects but struggle significantly with tasks requiring complex, abstract reasoning, such as spatial imagination, temporal ordering, and precise quantitative estimation. Overall, Gemini-2.5-Pro [team2023gemini] is the top one in the SIBench, while the GPT-5 [gpt5] leads in the SIBench-mini.
@@ A.3.1 Areas of Model Strength
In Table IV and Table V, models consistently perform well on tasks that rely on recognizing objects and simple relationships of objects. On the existence task, Qwen2.5-VL-72B [bai2025qwen2] achieves a perfect score of 1.0, and many others are above 0.9. This indicates that identifying whether an object is present in a scene is an easy task for current models. On the Spatial Relation task, InternVL-3-78B [zhu2025internvl3] leads at 0.8402 and Qwen2.5-VL-72B [bai2025qwen2] also receives 0.8216. This strength suggests that models have a solid grasp of static relational ability.
@@ A.3.2 Areas of Model Weakness
The models exhibit several key weaknesses, consistent with the analysis in Section 5.2.2. First, models are weak at some tasks of Foundational Perception, such as Object Size Estimation shown in Table IV and Table V. Second, models struggle with quantitative reasoning. This is evident in tasks Relative Distance in Table IV and Table V. The scores indicate that models like Gemini-2.5-Flash [team2023gemini] and InternVL-3-78B [zhu2025internvl3] rely on coarse visual clues rather than precise metric representations. Third, models show a near-total absence of spatial imagination. On the Spatial Imagination task, models like GPT-4o-mini [hurst2024gpt] and InternVL-2.5-78B-MPO [zhu2025internvl3] receive almost the lowest scores. This indicates that models fail on tasks that require the creation and manipulation of a mental imagination. Models cannot visualize and reason about objects or scenarios that are not explicitly depicted, which is a key component of human-like spatial intelligence. Second, models are weak at inferring spatial properties from different viewpoints. This is shown by the low scores in Multi-View Reasoning, Camera Pose, and Geometric Reasoning. Models like Qwen2.5-VL-7B [bai2025qwen2], LLaVA-OneVision-72B [li2024llava], Gemini-2.5-Flash [team2023gemini] cannot intuitively understand how a camera’s movement changes a scene’s perspective, leading to poor performance. Similarly, models struggle with temporal appearance order, with LLaVA-OneVision-7B [li2024llava] scoring 0.0000. This confirms that models’ unreliable ability for non-static scenes.
TABLE IV: Performance evaluation of different models on visual spatial reasoning tasks.
@@
Basic Perception Reach Prediction $0.6000$ 0.6750 0.6500 0.4750 0.6500 0.6000 0.5750
Height $0.5117$ 0.6117 0.5883 0.6167 0.6167 0.6517 0.5733
Existence 0.9750 0.9000 0.9250 0.9750 0.9750 0.9500 0.9000
Occlusion $0.5660$ 0.6620 0.5880 0.6160 0.6900 0.7940 0.6160
Object Shape $0.2308$ 0.2846 0.2692 0.3692 0.3154 0.3385 0.3000
Counting 0.5672 0.5042 0.4454 0.3991 0.4370 0.5462 0.4916
Object Size Estimation $0.5265$ 0.4679 0.5530 0.4575 0.5590 0.4573 0.6139
Spatial Understanding Spatial Compatibility $0.5280$ 0.4673 0.4813 0.4673 0.6168 0.6075 0.5748
Coordinate Conversion $0.5838$ 0.5413 0.5113 0.3563 0.6475 0.6475 0.5900
Trajectory Description $0.3205$ 0.5000 0.2179 0.4615 0.3718 0.2692 0.5513
Geometric Reasoning $0.2103$ 0.2738 0.2897 0.1786 0.2659 0.2857 0.2659
Spatial Imagination 0.3100 0.3075 0.2325 0.2575 0.2925 0.2800 0.3100
S Spatial Grid $0.4675$ 0.4025 0.4375 0.5650 0.6375 0.6425 0.7525
Temporal Appearance Order $0.1818$ 0.0000 0.2727 0.3636 0.2727 0.0909 0.2727
Multi-View Reasoning $0.3552$ 0.3892 0.3287 0.4358 0.4125 0.3934 0.3998
Situational QA $0.4994$ 0.4972 0.4729 0.4036 0.5329 0.5351 0.5173
Velocity Acceleration $0.3613$ 0.3882 0.2605 0.3344 0.4202 0.4840 0.5092
Relative Distance $0.4695$ 0.4629 0.3130 0.3608 0.5812 0.6324 0.5462
Camera Pose $0.4041$ 0.4477 0.3198 0.3459 0.5029 0.5581 0.5087
Spatial Relation 0.8239 0.7370 0.5892 0.4395 0.8130 0.8435 0.8133
Object Localization $0.4704$ 0.7290 0.5763 0.5670 0.3832 0.4922 0.2461
Planning Maze Navigation $0.3475$ 0.3375 0.2950 0.6400 0.5650 0.4775 0.4850
Route Planning $0.3220$ 0.3220 0.3729 0.4068 0.3220 0.2542 0.4576
Overall $0.4712$ 0.4850 0.4278 0.4389 0.5338 0.5481 0.5252
TABLE V: Performance evaluation of different models on the SIBench-mini.
@@
Basic Perception Reach Prediction $0.6000$ 0.6750 0.5750 0.6500 0.5000 0.5750 0.6000
Height $0.6500$ 0.6000 0.5633 0.7750 0.7500 0.7500 0.8000
Existence $0.9500$ 0.9500 1.000 0.9750 0.9500 0.9750 0.9250
Occlusion $0.6000$ 0.5500 0.6160 0.7250 0.8000 0.7250 0.6250
Object Shape $0.2500$ 0.2750 0.3000 0.4000 0.3250 0.2000 0.4750
Counting $0.8889$ 0.6111 0.4916 0.7222 0.9444 0.8333 0.7222
Object Size Estimation $0.5150$ 0.6675 0.6139 0.5800 0.5450 0.6175 0.6425
Spatial Understanding Spatial Compatibility $0.6250$ 0.7000 0.5748 0.7000 0.6500 0.7250 0.6250
Coordinate Conversion $0.7000$ 0.7000 0.5900 0.8000 0.7250 0.7000 0.6750
Trajectory Description 0.5500 0.6250 0.5513 0.3750 0.4500 0.4500 0.4750
Geometric Reasoning $0.2750$ 0.2500 0.2659 0.3000 0.3000 0.4000 0.3500
Spatial Imagination $0.3750$ 0.3000 0.3100 0.3250 0.3250 0.3250 0.3250
Spatial Grid $0.6250$ 0.8250 0.7265 0.9250 0.9750 1.0000 0.7750
Temporal Appearance Order $0.3636$ 0.5454 0.2727 0.1818 0.9091 0.1818 1.0000
Multi-View Reasoning $0.5250$ 0.3250 0.2269 0.3750 0.5250 0.4500 0.5750
Situational QA $0.3750$ 0.3438 0.5173 0.3438 0.5000 0.4375 0.6250
Velocity Acceleration $0.4000$ 0.4500 0.5092 0.4750 0.5750 0.5000 0.5250
Relative Distance $0.5325$ 0.5650 0.4811 0.6775 0.7350 0.6675 0.7075
Camera Pose 0.2250 0.4000 0.4884 0.3000 0.4000 0.5000 0.4000
Spatial Relation $0.6250$ 0.5750 0.8133 0.7500 0.6500 0.6750 0.6000
Object Localization $0.7250$ 0.7000 0.2430 0.7500 0.8250 0.7750 0.8000
Planning Maze Navigation $0.4000$ 0.5000 0.3150 0.7750 0.8500 0.7750 0.7500
Route Planning 0.3750 0.3000 0.4576 0.2500 0.4500 0.3750 0.4500
Overall $0.5252$ 0.5168 0.5355 0.5822 0.6216 0.6295 0.6906